# SimBEV: A Synthetic Multi-Task Multi-Sensor Driving Data Generation Tool and Dataset

### v1

@@ B.1 SimBEV configuration
We configured SimBEV toward generating diverse scenarios when creating the SimBEV dataset, and collected data from every sensor supported by SimBEV (RGB, semantic segmentation, instance segmentation, depth, and optical flow cameras; regular and semantic lidar; radar; GNSS; and IMU). The configuration used for each sensor type is listed in Tab. 6 and the arrangement of sensors on the ego vehicle is detailed in Tab. 7.
@@
Lidar 128 channels, 120.0 m range, 20.0 Hz rotation frequency, 5,242,880 points per second, 10.67 to -30.67 vertical FoV,
@@
Semantic lidar 128 channels, 120.0 m range, 20.0 Hz rotation frequency, 5,242,880 points per second, 10.67 to -30.67 vertical FoV
@@
Our sensor setup was mostly inspired by [2] (e.g. the 1600$\times$900 image resolution, the arrangement of the cameras, and the lidar’s vertical FoV), though there are a few differences. Our lidars (regular and semantic) have 128 channels instead of 32 to collect a much denser point cloud, which can be downsampled by the user later on if desired. For the GNSS and IMU we used bias and noise standard deviation values of a GNSS/INS module found in a typical experimental autonomous driving platform.
@@
SimBEV uses the probability distributions listed in Tab. 8 to randomize the parameters that control weather in CARLA. These distributions are interdependent to ensure the resulting weather is realistic (e.g. a combination of heavy rain and clear sky is unrealistic). Each of the configured parameters is briefly discussed below.
  • Cloudiness controls the amount of clouds. Values range from 0 to 100, 0 being a clear sky.
  • Precipitation controls the intensity of rain. Values range from 0 to 100, 0 being no rain.
  • Precipitation deposits controls the amount of puddles. Values range from 0 to 100, 0 being no puddles and 100 a road filled with water.
  • Wetness controls camera image blurriness caused by rain. Values range from 0 to 100, 0 being no blurriness.
  • Wind intensity controls the strength of wind. Values range from 0 to 100, 0 being no wind at all.
  • Sun azimuth angle controls the azimuth angle of the sun. Values range from 0 to 360.
  • Sun altitude angle controls the altitude angle of the sun. Values range from -90 to 90, -90 being midnight and 90 being midday.
  • Fog density controls fog concentration or thickness. Values range from 0 to 100, 0 being no fog.
  • Fog distance controls how far away the fog starts. Values can be any nonnegative number.
  • Fog falloff controls the density of the fog (as in specific mass). Values can be any nonnegative number. If set to 0, the fog will be lighter than air and will cover the whole scene. If set to 1, the fog is approximately as dense as air. For values greater than 5 the fog will be so dense that it will be compressed to the ground level. Fog falloff is set to 0.01 in Town12, Town13, and Town15 due to their non-zero elevation.
@@
Table 8: Probability distribution of SimBEV weather parameters. $\mathcal{B}$: beta distribution. $\mathcal{N}$: normal distribution. $\mathcal{U}$: uniform distribution. $\mathcal{LN}$: log-normal distribution.
@@
Distance between generated waypoints used for BEV ground truth calculation 0.4 m
Distance between generated waypoints used as spawning points 24.0 m
Number of spawned background vehicles ($s$: number of available spawn points) $\mathcal{U}_{i}(0,s-3)$
Number of spawned pedestrians $\mathcal{U}_{i}(0,640)$
@@
Probability of vehicles opening door(s) when stopped 10.0%
@@
Minimum pedestrian speed 0.8 m/s
Maximum pedestrian speed ($r$: minimum pedestrian speed) $\max(r,\mathcal{LN}(0.16,0.64))$ m/s
Minimum street light intensity 10,000 lm
Change in street light intensity ($m$: average street light intensity of the scene) $\mathcal{U}(-m,m)$ lm
@@
B.2 SimBEV dataset statistics
The SimBEV dataset comprises 102,400 annotated frames, 8,315,935 3D bounding boxes (3,792,499 of which are valid), and 2,793,491,357 BEV ground truth labels, broken down by class in Tab. 10. Cars and pedestrians make up the largest share of 3D bounding boxes, though they include a large number of motorcycles and bicycles as well. This makes sense given that the majority of models in CARLA’s vehicle library are cars (as opposed to e.g. only one bus model). BEV labels are dominated by the road class, followed by cars, trucks, and buses due to their larger footprint compared to the rest.
@@ B.2 SimBEV dataset statistics
Table 10: Breakdown of total and valid 3D bounding boxes and BEV ground truth labels by class for the SimBEV dataset.
As discussed previously, SimBEV randomizes CARLA’s weather parameters according to the distributions specified in Tab. 8. Figure 8 shows the distribution of weather across the SimBEV dataset, where precipitation (rain intensity, $k_{p}$) and fog density ($k_{f}$) values for each scene are categorized into none ($<$10%), low (10 - 40%), moderate (40 - 70%), and heavy (70 - 100%); while sun altitude angle ($k_{al}$) is categorized into night (-90 - 0 deg), dawn/dusk (0 - 6 deg), and day (6 - 90 deg). Figure 8 shows that SimBEV contains a good mix of different weather conditions, with either rain or fog present in around half of the scenes and nearly a quarter of the scenes occurring at night.
@@
Figure 9: Distribution of the number of spawned vehicles (includes cars, trucks, buses, motorcycles, and bicycles) and pedestrians across SimBEV dataset scenes.
@@
Looking at the distribution of the number of spawned vehicles (includes cars, trucks, buses, motorcycles, bicycles) and pedestrians among SimBEV dataset scenes shown in Fig. 9, it is clear that the scenes range from relatively empty to congested and crowded. The distribution of pedestrians is supposed to be uniform, bus CARLA often spawns fewer pedestrians than requested, and the number of unspawned pedestrians grows rapidly when hundreds of pedestrians are requested. Hence there are many scenes with 240 - 320 pedestrians and very few with more than 480.
@@
Figure 14: Logarithmic BEV heat map of the SimBEV dataset for different classes.
@@
Because Fig. 9 shows the total number of spawned vehicles and pedestrians, many of which may be far from the ego vehicle, it does not paint the full picture of what the ego vehicle observes. Hence in Fig. 10 we break down the number of valid 3D object bounding boxes per frame by class for the SimBEV dataset. The distribution of the bounding boxes is similar to [2], though our dataset offers a sizable number of frames with many (65+) valid car or pedestrian bounding boxes as well. As expected, due to fewer models in CARLA’s vehicle library, the vast majority of frames only include a handful of trucks, buses, motorcycles, and bicycles.
@@
For more insight into the SimBEV dataset we turn to Fig. 11 through Fig. 13. Figure 11 shows that 3D object bounding box distances from the ego vehicle are nearly uniformly distributed for all classes, in contrast to [2], which is likely due to the higher density and range of our lidar point cloud. Figure 12 shows a reasonable speed range for all object classes, which is comparable to [2] with a few exceptions. We have a large number of running pedestrians (3+ m/s) which can act as edge cases for perception and behavior prediction algorithms. For other classes, our data is collected from both urban and highway environments (unlike [2] which only collected data from urban environments), leading to many fast-moving objects. Figure 13 shows the distribution of the number of lidar and radar points within valid 3D object bounding boxes with respect to distance. Consistent with [2], larger object bounding boxes have more points inside and the number of points for all classes decrease with increasing distance.

### v2

@@ B.1 SimBEV configuration
We configured SimBEV to generate a diverse set of unique scenarios for the SimBEV dataset, and collected data from all sensor types supported by SimBEV (RGB, semantic segmentation, instance segmentation, depth, and optical flow cameras; regular and semantic lidar; radar; GNSS; and IMU). Sensor configurations are listed in Tab. 6 and the arrangement of the sensors on the ego vehicle is shown in Fig. 9 and Fig. 10, and detailed in Tab. 7.
@@
Lidar 128 channels, 120.0 m range, 20.0 Hz rotation frequency, 5,242,880 points per second, -30.67 to 10.67 vertical FoV,
@@
Semantic lidar 128 channels, 120.0 m range, 20.0 Hz rotation frequency, 5,242,880 points per second, -30.67 to 10.67 vertical FoV
@@
Our sensor setup was mostly inspired by [3] (e.g. the 1600$\times$900 image resolution, the arrangement of the cameras, and the lidar’s vertical FoV), though there are a few differences. Our lidars (both regular and semantic) have 128 channels instead of 32 to collect a much denser point cloud, which can be downsampled by the user later on if desired. For GNSS and IMU, we used the bias and noise standard deviation values of a GNSS/INS module found in a typical experimental autonomous driving platform.
@@
Table 8: Probability distribution used in SimBEV for weather parameters. $\mathcal{B}$: beta distribution. $\mathcal{N}$: normal distribution. $\mathcal{U}$: uniform distribution. $\mathcal{LN}$: log-normal distribution.
SimBEV uses the probability distributions listed in Tab. 8 to randomize the parameters that control the weather in CARLA. These distributions are interdependent to ensure that the resulting weather is realistic (e.g. a combination of heavy rain and clear sky is unrealistic). Each of the configured parameters is briefly discussed below.
Figure 10: Coordinate frames of the perception sensors used in SimBEV to create the SimBEV dataset.
@@
Distance between CARLA-generated waypoints used for BEV ground truth calculation 0.4 m
Distance between CARLA-generated waypoints used as vehicle spawn location 24.0 m
Number of background vehicles ($s$: number of available spawn locations) $\mathcal{U}_{i}(0,s-3)$
Number of pedestrians $\mathcal{U}_{i}(0,640)$
@@
Probability of vehicle door(s) getting open when stopped 10.0%
@@
Minimum speed of pedestrians 0.8 m/s
Maximum speed of pedestrians ($r$: minimum pedestrian speed) $\max(r,\mathcal{LN}(0.16,0.64))$ m/s
Minimum intensity of street lights 10,000 lm
Change in the intensity of street lights ($m$: average intensity of all street lights in the scene) $\mathcal{U}(-m,m)$ lm
@@
Table 10: Breakdown of the number of total and valid 3D bounding boxes and BEV ground truth labels by class for the SimBEV dataset.
@@
  • Cloudiness ($k_{c}$) controls the volume of clouds. Values range from 0 to 100.
  • Precipitation ($k_{p}$) controls the intensity of rain. Values range from 0 to 100.
  • Precipitation deposits ($k_{pd}$) controls the amount of puddles. Values range from 0 to 100, with 0 being no puddles and 100 a road filled with water.
  • Wetness ($k_{w}$) controls the intensity of camera image blurriness caused by rain. Values range from 0 to 100.
  • Wind intensity ($k_{wi}$) controls the strength of wind. Values range from 0 to 100.
  • Sun azimuth angle ($k_{az}$) controls the azimuth angle of the sun. Values range from 0 to 360.
  • Sun altitude angle ($k_{al}$) controls the altitude angle of the sun. Values range from -90 to 90, with -90 representing midnight and 90 midday.
  • Fog density ($k_{f}$) controls fog concentration or thickness. Values range from 0 to 100, with 0 being no fog.
  • Fog distance ($k_{fd}$) controls how far away the fog starts, and can be any nonnegative number.
  • Fog falloff ($k_{ff}$) controls the density of the fog (as in specific mass), and can be any nonnegative number. If set to 0, the fog will be lighter than air and will cover the whole scene. If set to 1, the fog is approximately as dense as air. For values greater than 5 the fog will be so dense that it will be compressed to the ground level. Fog falloff is set to 0.01 in Town12, Town13, and Town15 due to their non-zero elevation.
B.2 SimBEV dataset statistics
The SimBEV dataset comprises 102,400 annotated frames, 8,315,935 3D object bounding boxes (3,792,499 of which are valid), and 2,793,491,357 BEV ground truth labels, broken down by class in Tab. 10. Cars and pedestrians make up the largest share of 3D object bounding boxes, though those boxes include a large number of motorcycles and bicycles as well. This makes sense since the majority of models in CARLA’s vehicle library are cars (compared to, e.g., only one bus model). BEV labels are dominated by the road class, followed by car, truck, and bus due to their larger footprint compared to the rest.
As discussed previously, SimBEV randomizes CARLA’s weather parameters according to the distributions specified in Tab. 8. Figure 11 shows the distribution of weather across the SimBEV dataset, where precipitation (rain intensity, $k_{p}$) and fog density ($k_{f}$) values for each scene are categorized into none ($<$10%), low (10 - 40%), moderate (40 - 70%), and heavy (70 - 100%); while sun altitude angle ($k_{al}$) is categorized into night (-90 - 0 deg), dawn/dusk (0 - 6 deg), and day (6 - 90 deg). Figure 11 shows that SimBEV contains a good mix of different weather conditions, with rain or fog present in about half of the scenes and nearly a quarter of the scenes occurring at night.
Looking at the distribution of the number of spawned vehicles (cars, trucks, buses, motorcycles, bicycles) and pedestrians across scenes of the SimBEV dataset, shown in Fig. 12, it is clear that the scenes range from relatively empty to congested and crowded. The distribution of pedestrians is supposed to be uniform, but CARLA often spawns fewer pedestrians than requested, and the number of unspawned pedestrians grows rapidly when hundreds of pedestrians are requested. Moreover, in some cases CARLA cannot spawn pedestrians because there are no walkable areas around the ego vehicle (e.g., when the ego vehicle is traveling on a rural road). Hence, there are many scenes with 0 and 240 - 320 pedestrians and very few with more than 480.
@@ B.2 SimBEV dataset statistics
Figure 12: Distribution of the number of spawned vehicles (cars, trucks, buses, motorcycles, and bicycles) and pedestrians across the scenes of the SimBEV dataset.
@@
Figure 17: Logarithmic BEV heat maps of the SimBEV dataset for different classes.
@@
Because Fig. 12 shows the total number of spawned vehicles and pedestrians, many of which may be far from the ego vehicle, it may not fully represent what the ego vehicle observes. Hence in Fig. 13 we break down the number of valid 3D object bounding boxes per frame by class. The distribution of the bounding boxes is similar to [3], although our dataset offers a sizable number of frames with many (65+) valid car/pedestrian bounding boxes as well. As expected, due to having fewer models in CARLA’s vehicle library, the vast majority of frames only include a handful of trucks, buses, motorcycles, and bicycles.
@@
Figure 14 to Fig. 16 provide more insight into the SimBEV dataset. Figure 14 shows that the distances of valid 3D object bounding boxes from the ego vehicle are nearly uniformly distributed for all classes, in contrast to [3], which is likely due to the higher density and range of our lidar point cloud. Figure 15 shows a reasonable speed range for all classes, which is comparable to [3] with a few exceptions. Our dataset has a large number of running pedestrians (3+ m/s), which can serve as edge cases for perception and behavior prediction algorithms. For other classes, our data was collected from both urban and highway environments (unlike [3], which only collected data from urban environments), leading to many fast-moving objects. Fig. 16 shows the distribution of the number of lidar and radar points within valid 3D object bounding boxes with respect to distance from the ego vehicle. Consistent with [3], larger object bounding boxes have more points inside and the number of points for all classes decreases with increasing distance.